OpenStudy (anonymous):

I am in terrible need of help here, like i don't get this at all and it NEEDS to be done

OpenStudy (anonymous):

The function H(t) = −16t2 + 90t + 50 shows the height H(t), in feet, of a projectile after t seconds. A second object moves in the air along a path represented by g(t) = 28 + 48.8t, where g(t) is the height, in feet, of the object from the ground at time t seconds. Part A: Create a table using integers 1 through 4 for the 2 functions. Between what 2 seconds is the solution to H(t) = g(t) located? How do you know? (6 points) Part B: Explain what the solution from Part A means in the context of the problem. (4 points)

OpenStudy (unklerhaukus):

t H(t) = −16t2 + 90t + 50 1 H(1) = -16(1)^2+90(1)+50 = -16+90+50 = 124 2 H(2) = -16(2)^2+90(2)+50 = = 3 4

OpenStudy (usukidoll):

part a. let t = 1,2,3,4 and plug it into the 2 functions.

OpenStudy (anonymous):

What about our first reply here, is that any good?

OpenStudy (anonymous):

not... woop woop

OpenStudy (usukidoll):

all I see are mods deleting stuff. . . . other than that idk. but part a seems to have been partly started by Unkle.

OpenStudy (usukidoll):

we need to find h(t) = g(t) like at what time does those functions equal to each other.

OpenStudy (usukidoll):

boo. I don't like these explain questions like in part b. LIke I can see through part A, but why a short essay answer on B?

OpenStudy (anonymous):

as long as we can conjure up a solution to a i might be able to concoct something for B

OpenStudy (anonymous):

we just really need to finish this. there is so much on my shoulders now and i just want this out of the way